# Multipartite entanglement distribution in a topological photonic network

### v1

@@ Abstract
In the ongoing effort towards a scalable quantum computer, multiple technologies have been proposed. Some of them exploit topological materials to process quantum information. In this work, we propose a lattice of photonic cavities with alternating hoppings to create a modified multidomain SSH chain, that is, a sequence of topological insulators made from chains of dimers. A qubit is then coupled to each boundary. We show this system is well suited for quantum information processing because topological transfer of photons through this one-dimensional lattice can entangle any set of qubits on demand, providing a scalable quantum platform. We verify this claim evaluating entanglement measures and witnesses proving that bipartite and multipartite entanglement is produced, even in the presence of some disorder.
I Introduction
Quantum mechanics allows for the creation of a computer able to solve problems with a higher performance than classical ones Feynman (1982); Shor (1994); Grover (1996). Plenty of research nowadays deals with this statement, from the study of algorithms and metrics that prove quantum advantage Boixo et al. (2018); Bravyi et al. (2018) to the design of hardware primitives able to create said computer Cirac and Zoller (1995); Zhong et al. (2020). In this work we focus on the latter, the physical elements of a quantum computer, and analyse them with a metric that is a necessary condition to quantum advantage: multipartite entanglement production Nielsen (2006). A list of proposals for quantum computers includes superconducting circuits Blais et al. (2021), ion or cold atoms traps Pogorelov et al. (2021), photonic systems, quantum dots, and topological materials, to name but a few. It is unclear which platform will have the most impact, but the history of classical computing teaches us that multiple implementations coexisted and all are worth researching. Here we focus on a topological photonic lattice coupled to color center emitters, given the outstanding progress taking place on those respective areas of research, that we briefly introduce now.
@@ I Introduction
Topological phases of matter are an interesting quantum platform because of their robustness against some kinds of disorder. This protection helps with the survival of quantum properties for longer times, the main limiting factor in quantum computing applications. One of the applications topological matter finds in quantum technologies is the use of 1D symmetry-protected edge states to transfer information across the material Bello et al. (2016, 2017); Lang and Büchler (2017); Longhi et al. (2019); Longhi (2019); Yuce (2019); Pérez-González et al. (2019); D’Angelis et al. (2020). This transfer can take place along several regions with different topological characteristics, that is, along multiple domains. Recently, it was shown that a multidomain approach drastically accelerates transfer Zurita et al. (2023), which greatly increases the technological potential of this phenomenon. In that regard, photonic lattices are experiencing an outstanding growth in the number and complexity of topological models they can implement Malkova et al. (2009); Wang et al. (2009); Hafezi et al. (2013); Weimann et al. (2017); St-Jean et al. (2017); Zhao et al. (2018); Yang et al. (2018); Huda et al. (2020); Zurita et al. (2020); Gong et al. (2021); Pernet et al. (2022); Rosiek et al. (2023); Kim et al. (2022) and use in applications like lasing Bandres et al. (2018); Contractor et al. (2022), sensing Parto et al. (2023) or, as mentioned, quantum information processing Blanco-Redondo et al. (2018); Hu et al. (2020); Tschernig et al. (2021); Dai et al. (2022). In particular, multidomain SSH-like models have been implemented experimentally in several platforms Blanco-Redondo et al. (2016); Zhao et al. (2018); Huda et al. (2020); Pernet et al. (2022).
@@
Another reason to turn our attention to photonics is their ability to couple to emitters, forming highly scalable platforms that are created using advanced fabrication techniques Khoury and Abbarchi (2022). Recent breakthroughs have been achieved like the production of high-brightness G-centers and the control of their position Hollenbach et al. (2020, 2022), the coupling of G-centers to a photonic cavity in a single-photon basis Baron et al. (2022); Redjem et al. (2023), or the creation of vacancies or W-, T- and G-centers coupled to waveguides Tait et al. (2020); MacQuarrie et al. (2021); DeAbreu et al. (2022); Prabhu et al. (2023); Lukin et al. (2023).
@@
The advancements cited so far have led to theoretical works about whether these emitters can be used as qubits of a quantum computer, coupled to each other through the edge states of a photonic topological system Lang and Büchler (2017); Yan et al. (2021); Hollenbach et al. (2020); Guimbao et al. (2021). In this work we address this question, that is, we study whether coherent operation of multi-qubit states is possible. Furthermore, we propose protocols for coherent operation that scale well with the number of qubits. To show this we propose a nearly one-dimensional system and prove that it produces entanglement between any pair or bigger subset of qubits in experimentally feasible settings, proving that coherent processes mediated by the edge states overcome dissipation and decoherence.
@@
The structure of this work is as follows: in Section II, we introduce the system, consisting of a quasi-1D topological photonic lattice, with qubits placed at its interfaces. We also propose an experimental implementation for it, using color center qubits in a silicon cavity lattice. Next, we demonstrate in Section III how the topological transfer of a photon through the lattice can be used to prepare a Bell state between arbitrary pairs of these qubits. To quantify the performance of the protocol we study the concurrence between the target qubits in the presence of different kinds of disorder. We then use similar topological transfers in Section IV to prepare the W and GHZ states between three qubits and evaluate the performance of the protocol using genuine multipartite entanglement witnesses in the same parameter regimes as with Bell states. Finally, we state our conclusions about the usability and scalability of this system as a quantum computing platform, together with possible future lines of research.
II Description of the system
@@
II.1 Topological photonic lattice
@@
II.2 Coupled qubits
Qubits can be coupled to any number of extremal cavities, as shown in Fig. 1 (a,b), and we will label them starting at 0. They will act like photon emitters with the frequency of the cavity. In one of the protocols, we will extend the qubits to qutrits, so we keep the notation general. The coupling between the qubits and the cavities housing them is, in the rotating wave approximation:
@@ II.2 Coupled qubits
where $j(k),\alpha(k)$ labels the cavity where the $k$-th qudit is located, $c_{j(k),\alpha(k)}$ destroys a photon in that cavity, $p$ is the number of qudits, and $|s_{l}\rangle$ and $|s_{u}\rangle$ are the lower and upper qudit states connected by the cavity-mediated transition, with $s_{l,u}=0,\ldots,d-1$, where $d$ is the dimension of the qudits used.
The odd domains in this model are SSH chains, while the even ones can be obtained as a limiting case of the extended SSH model studied in Pérez-González et al. (2019). The use of domains with topological invariants of ±$\pm 1$ (see Appendix A) provides domain walls with two states each, one of them can be used to interface with a qubit while the other is used to transfer information along the wall. This allows us to create a 1D network of qubits with full connectivity. A system with $N$ domains then has $2N$ boundary states, that are topologically protected against off-diagonal disorder as long as $w$ is greater than all $v$-type (intra-domain) and $u$-type (stub) links, as shown in Fig. 2 (a,b). Some of these states are mainly localized in a single cavity and can thus be easily coupled to a qubit (the left and right states $|\mathcal{L}\rangle$, $|\mathcal{R}\rangle$ and the $|\mathcal{S}_{k}\rangle$ states inside the domain walls), while others ($|\mathcal{P}_{k}\rangle$) can be used to transfer information along a wall. These last states disappear if we set $v_{k}>w$ in one of their contiguous domains. We show this, as well as the extent of the topological protection, in Appendix A.
All of the edge states that are not decoupled from the rest of the system (e.g. by setting their $u_{k}=0$) will hybridize, forming bonding and antibonding eigenstates. The topological protection ensures that they will stay decoupled from all bulk states. The effective overlap between each pair of states depends on the links between them, given that these dictate their decay length.
Since the effective overlap between each pair of states depends on the links between them, a transfer protocol between the two ends of all the hybridized states can be designed by tuning some of the hopping amplitudes. Thus, by tuning the weaker hopping amplitudes, marked in orange in Fig. 2 (a), the dynamics in this subspace can be controlled. Using the right pulses, a photon initially on the leftmost state of the set of hybridized states can always be made to end up on the rightmost one, as explored in Zurita et al. (2023). We show the pulses we used in Appendix B. This transfer process, which we will refer to as $T$, is always topologically protected and can be established between any two domain boundaries.
@@
Apart from connectivity, there is an additional advantage to using domains. As discussed in the literature Lang and Büchler (2017); Zurita et al. (2023), a downside of bidirectional transfer protocols in 1D topological insulators is the exponential scaling of their duration as a function of distance, which in turn allows small symmetry-breaking perturbations to become catastrophic for the performance of the protocol. This is caused by their exponential decay into the bulk, which determines their effective overlap. This problem can be solved by adding intermediate domain walls between the ends of the model, as shown in Zurita et al. (2023). This makes the transfer duration grow only linearly with distance.
II.3 Experimental proposal
@@
The cavity lattice can be built using a triangular lattice of holes in a slab of silicon, where the absence of holes creates a cavity, as shown in Fig. 1 (a). The hopping amplitude between contiguous cavities can be modified by changing the medium between cavities, by driving the cavity frequencies Liao et al. (2010) or using ancilliary cavities (either containing impurities Qin and Nori (2016) or with high frequencies Liao et al. (2010); Mukherjee et al. (2018)), or even by implementing nanowire-induced movable cavities, as shown experimentally in Birowosuto et al. (2014). As we explain below, the ability to tune these amplitudes is not necessary to implement the two-qubit protocol, although it does increase the resulting fidelity.
@@
The qubits can be impurities in the silicon like G- or T-centers, located within the appropriate cavities, as shown in Fig. 1 (b). The qudit-cavity coupling parameters, $g_{k}$, could be modified by tuning the energies of the impurity with an electrical gate in or out of resonance with the cavity Hollenbach et al. (2020), by tuning the cavity frequency as mentioned above, or otherwise by using an indirect transition with an auxilliary laser Lang and Büchler (2017).
As mentioned in Hollenbach et al. (2020); Redjem et al. (2023), controlling the exact placement and orientation of the impurities can pose a challenge, although advances are being made with controllable growth techniques Hollenbach et al. (2020, 2022); Redjem et al. (2023). A solution could be to create the cavity lattice around the impurities after they are placed Redjem et al. (2023). The topological robustness against off-diagonal disorder in our model greatly dampens the effects of the fabrication defects that these systems may have, especially the distances between cavities.
Figure 2: (a) Photonic cavity lattice with $N=4$ domains and $\ell=4$ cavities per domain. Qubits (in green) are coupled to the cavities at the ends and domain walls, with a coupling amplitude of $g_{k}$. Hopping amplitudes used as control parameters are shown in orange: we use $u$ for extremal links, $v$ for all the rest. (b) Topologically protected states in this system. Two chiral states, $\mathcal{S}$ and $\mathcal{P}$, are located at each wall. States in odd (even) boundaries are completely located in sublattice $B$ ($A$), and shown in blue (yellow). (c) Transfer process between the left end and the first wall. The parameter $v_{\mathrm{tr}}<w$ allows the transfer, while $v_{\mathrm{bar}}\gg w$ stops it. (d) Transfer process between the two ends. The $\mathcal{P}$ states are involved now. Two different pulses are used for the extremal ($v_{\mathrm{tr}}^{(1)}$) and central ($v_{\mathrm{tr}}^{(2)}$) domains to achieve a high-fidelity transfer.
III Bell state preparation
III.1 Preparation protocol
@@
In this section, we use this photonic multidomain topological system to prepare maximally entangled states on two qubits, i.e. Bell states. In order to explain the protocol, we need to define some notation. Let $n_{d}$ be the total number of domains and $\ell$ the number of sites per domain. As explained above, qubits sit on the stubs between domains. Let $R$ and $L$ stand for the rightmost and leftmost qubits, which are the furthest apart and so the hardest to entangle. Therefore, we describe the protocol for these qubits only, given that the protocol for any other pair is analogous. The ground and excited states of qubit $L$ are |$\ket{0}_{L}$ and |$\ket{1}_{L}$, repectively, and we will refer to its coupling to the photonic cavity stub here as $g_{L}$. Furthermore, since coherent single-photon emission has been reported experimentally for color-center qubits Baron et al. (2022); Redjem et al. (2023), we consider $\pi$- and $\pi/2$-pulses to be an accesible operation between qubit $L$ and its cavity  [^1] We use sine-shaped pulses for all $g$ parameters, see Appendix B for details.. We label the latter unitary as $(\Pi/2)_{01}^{(L)}$. We use similar notation for qubit $R$. The vacuum state of the photonic lattice is |$\ket{\varnothing}_{\gamma}$. The topological transfer of a photonic excitation from the leftmost cavity to the rightmost through the edge states is $T_{LR}$.
@@
|$\ket{\Phi}_{LR}=\frac{1}{\sqrt{2}}\left[\ket{10}+e^{i\theta}\ket{01}\right]$ (6)
@@
$e^{i\theta}=[(-1)^{\ell/2+1}i]^{n_{d}},$ (7)
See Appendix C for a detailed derivation of this relative phase  [^2] This formula is related to those given in Zurita et al. (2023) for the SSH chain, where $\ell$ was defined differently and $n_{d}=n_{w}+1$..
III.2 Performance against disorder: concurrence
While the calculations reported so far demonstrate that this system generates entanglement in the ideal case, we need to explore the impact of disorder if we are to consider the system a scalable quantum platform. First of all, the final state will no longer be Eq. 6. Moreover, the final state has a nonzero probability amplitude of having a leftover photon in the photonic lattice, leading to decoherence in the state of the qubits. In this case, entanglement certification is not as simple as inspecting the state in Eq. 6, so instead we turn to concurrence, an entanglement measure Hill and Wootters (1997), to perform this task. We remind the reader that, if a two-qubit state has high concurrence, any latter state with lower concurrence may be prepared from the former with only local operations and classical communication (LOCC); in particular, without using the photonic lattice. In this sense, Bell states have maximal concurrence. We now turn our attention to the disorder model used, and then show that the protocol can produce high-concurrence states despite disorder.

### v2

@@ Abstract
In the ongoing effort towards a scalable quantum computer, multiple technologies have been proposed. Some of them exploit topological materials to process quantum information. In this work, we propose a lattice of photonic cavities with alternating hoppings to create a modified multidomain SSH chain, that is, a sequence of topological insulators made from chains of dimers. A qubit is then coupled to each boundary. We show this system is well suited for quantum information processing because topological transfer of photons through this one-dimensional lattice can entangle any set of qubits on demand, providing a scalable quantum platform. We verify this claim evaluating entanglement measures and witnesses, proving that bipartite and multipartite entanglement is produced, even in the presence of some disorder.
1 Introduction
Quantum mechanics allows for the creation of a computer able to solve problems with a higher performance than classical ones [1, 2, 3]. Plenty of research nowadays deals with this statement, from the study of algorithms and metrics that prove quantum advantage [4, 5], to the design of hardware primitives able to create said computer [6, 7]. In this work we focus on the latter, the physical elements of a quantum computer, and analyse them with a metric that is a necessary condition to quantum advantage: multipartite entanglement production [8]. A list of proposals for quantum computers includes superconducting circuits [9], ion or cold atoms traps [10], photonic systems, quantum dots, and topological materials, to name but a few. It is unclear which platform will have the most impact, but the history of classical computing teaches us that multiple implementations coexisted and all are worth researching. Here we focus on a topological photonic lattice coupled to color center emitters, given the outstanding progress taking place on those respective areas of research, that we briefly introduce now.
@@ 1 Introduction
Topological phases of matter are an interesting quantum platform because of their robustness against some kinds of disorder. This protection helps with the survival of quantum properties for longer times, the main limiting factor in quantum computing applications. One of the applications topological matter finds in quantum technologies is the use of 1D symmetry-protected edge states to transfer information across the material [11, 12, 13, 14, 15, 16, 17, 18]. This transfer can take place along several regions with different topological characteristics, that is, along multiple domains. Recently, it was shown that a multidomain approach drastically accelerates transfer [19], which greatly increases the technological potential of this phenomenon. In that regard, photonic lattices are experiencing an outstanding growth in the number and complexity of topological models they can implement [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and use in applications like lasing [34, 35], sensing [36] or, as mentioned, quantum information processing [37, 38, 39, 40]. In particular, multidomain SSH-like models have been implemented experimentally in several platforms [41, 25, 27, 31].
@@
Another reason to turn our attention to photonics is their ability to couple to emitters, forming highly scalable platforms that are created using advanced fabrication techniques [42]. Recent breakthroughs have been achieved like the production of high-brightness G-centers and the control of their position [43, 44], the coupling of G-centers to a photonic cavity in a single-photon basis [45, 46], and the creation of vacancies or W-, T- and G-centers coupled to waveguides [47, 48, 49, 50, 51].
@@
The advances cited so far have led to theoretical works about whether these emitters can be used as qubits of a quantum computer, coupled to each other through the edge states of a photonic topological system [13, 52, 43, 29, 53]. In this work we address this question, that is, we study whether coherent operation of multi-qubit states is possible. Furthermore, we propose protocols for coherent operation that scale well with the number of qubits. To show this we propose a nearly one-dimensional system, and prove that it produces entanglement between any pair or bigger subset of qubits in experimentally feasible settings, proving that coherent processes mediated by the edge states overcome dissipation and decoherence.
@@
The structure of this work is as follows: in Section 2, we introduce the system, consisting of a quasi-1D topological photonic lattice, with qubits placed at its interfaces. We also propose an experimental implementation for it, using color center qubits in a silicon cavity lattice. Next, we demonstrate in Section 3 how the topological transfer of a photon through the lattice can be used to prepare a Bell state between arbitrary pairs of these qubits. To quantify the performance of the protocol we study the concurrence between the target qubits in the presence of different kinds of disorder. We then use similar topological transfers in Section 4 to prepare the W and GHZ states between three qubits and evaluate the performance of the protocol using genuine multipartite entanglement witnesses in the same parameter regimes as with Bell states. In Section 5, we explore the main sources of decoherence and loss affecting our protocols. Finally, we state our conclusions about the usability and scalability of this system as a quantum computing platform, together with possible future lines of research.
2 Description of the system
@@
2.1 Topological photonic lattice
@@
Figure 2: (a) Photonic cavity lattice with $N=4$ domains and $\ell=4$ cavities per domain. Qubits (green) are coupled to the ends and domain walls, with an amplitude of $g_{k}$. The links used as control parameters are shown in orange: we use $u$ for extremal links, $v$ for the rest. (b) Topologically protected states in this system. Two chiral states, $\mathcal{S}$ and $\mathcal{P}$, exist at each wall. States in odd (even) boundaries are confined to sublattice $B$ ($A$), and shown in blue (yellow). (c) Transfer between the left end and the first wall. The parameter $v_{\mathrm{tr}}<w$ allows the transfer, while $v_{\mathrm{bar}}\gg w$ stops it. (d) End-to-end transfer. The $\mathcal{P}$ states are involved now. Two different pulses are used for the extremal ($v_{\mathrm{tr}}^{(1)}$) and central ($v_{\mathrm{tr}}^{(2)}$) domains to achieve a high-fidelity transfer.
2.2 Coupled qubits
Qubits can be coupled to any number of extremal cavities, as shown in Fig. 1 (a,b), and we will label them starting at 0. They will act like photon emitters with the frequency of the cavity. In one of the protocols, we will extend the qubits to qutrits, so we keep the notation general. The coupling between the qudits and the cavities housing them is, in the rotating wave approximation:
@@ 2.2 Coupled qubits
where $j(k),\alpha(k)$ labels the cavity where the $k$-th qudit is located, $c_{j(k),\alpha(k)}$ destroys a photon in that cavity, $p$ is the number of qudits, $g_{k}$ is the coupling coefficient for the $k$-th qudit, and $|s_{l}\rangle$ and $|s_{u}\rangle$ are the lower and upper qudit states connected by the cavity-mediated transition, with $s_{l,u}=0,\ldots,d-1$, where $d$ is the dimension of the qudits used.
The odd domains in this model are SSH chains, while the even ones can be obtained as a limiting case of the extended SSH model studied in [17]. The use of domains with topological invariants of ±$\pm 1$ (see Appendix A) provides domain walls with two states each, one of them can be used to interface with a qubit while the other is used to transfer information along the wall. This allows us to create a 1D network of qubits with full connectivity. A system with $N$ domains then has $2N$ boundary states, that are topologically protected against off-diagonal disorder as long as $w$ is greater than all $v$-type (intra-domain) and $u$-type (stub) links, as shown in Fig. 2 (a,b). Some of these states are mainly localized in a single cavity and can thus be easily coupled to a qubit (the left and right states $|\mathcal{L}\rangle$, $|\mathcal{R}\rangle$ and the $|\mathcal{S}_{k}\rangle$ states inside the domain walls), while others ($|\mathcal{P}_{k}\rangle$) can be used to transfer information along a wall. These last states disappear if we set $v_{k}>w$ in one of their contiguous domains. We show this, as well as the extent of the topological protection, in Appendix A.
All edge states that are not decoupled from the rest of the system (e.g. by setting their $u_{k}=0$) will hybridize, forming bonding and antibonding eigenstates. The topological protection ensures that they will stay decoupled from all bulk states. The effective overlap between each pair of states depends on the links between them, given that these dictate their decay length.
For this reason, a transfer protocol between the two ends of all the hybridized states can be designed by tuning some of the hopping amplitudes. In fact, by tuning only the weaker hopping amplitudes, marked in orange in Fig. 2 (a), the dynamics in this subspace can be fully controlled. Using the right pulses, a photon initially on the leftmost state of the set of hybridized states can always be made to end up on the rightmost one, as explored in [19]. The effective Hamiltonian of the protected subspace in the SSSH chain is analogous to that of the Creutz ladder in the mentioned work. We show the pulses we used in Appendix B. This transfer process, which we will refer to as $T$, is always topologically protected and can be established between any two domain boundaries.
@@
Apart from connectivity, there is an additional advantage to using domains. As discussed in the literature [13, 19], a downside of bidirectional transfer protocols in 1D topological insulators is the exponential scaling of their duration as a function of distance, which in turn allows small symmetry-breaking perturbations to become catastrophic for the performance of the protocol. This is caused by their exponential decay into the bulk, which determines their effective overlap. This problem can be solved by adding intermediate domain walls between the ends of the model, as shown in [19]. This makes the transfer duration grow only linearly with distance.
2.3 Experimental proposal
@@
The cavity lattice can be built using a triangular lattice of holes in a slab of silicon, where the absence of holes creates a cavity, as shown in Fig. 1 (a). The hopping amplitude between contiguous cavities can be modified by changing the medium between cavities, by driving the cavity frequencies [54] or using ancillary cavities (either containing impurities [55] or with high frequencies [54, 56]), or even by implementing nanowire-induced movable cavities, as shown experimentally in [57]. As we explain below, the ability to tune these amplitudes is not necessary to implement the two-qubit protocol, although it does increase the resulting fidelity.
@@
The qubits can be impurities in the silicon like G- or T-centers, located within the appropriate cavities, as shown in Fig. 1 (b). The qudit-cavity coupling parameters, $g_{k}$, could be modified by tuning the energies of the impurity with an electrical gate in or out of resonance with the cavity [43], by tuning the cavity frequency as mentioned above, or otherwise by using an indirect transition with an auxiliary laser [13].
In Section 5, we discuss the main challenges our protocols would face in this platform, and how they could be overcome or mitigated using the topological features of the lattice.
3 Bell state preparation
3.1 Preparation protocol
@@
In this section, we use this photonic multidomain topological system to prepare maximally entangled states on two qubits, i.e. Bell states. In order to explain the protocol, we need to define some notation. Let $n_{d}$ be the total number of domains and $\ell$ the number of sites per domain. As explained above, qubits sit on the stubs between domains. Let $R$ and $L$ stand for the rightmost and leftmost qubits, which are the furthest apart and so the hardest to entangle. Therefore, we describe the protocol for these qubits only, given that the protocol for any other pair is analogous. The ground and excited states of qubit $L$ are |$\ket{0}_{L}$ and |$\ket{1}_{L}$, respectively, and we will refer to its coupling to the photonic cavity stub here as $g_{L}$. Furthermore, since coherent single-photon emission has been reported experimentally for color-center qubits [45, 46], we consider $\pi$- and $\pi/2$-pulses to be an accessible operation between qubit $L$ and its cavity [^1] We use sine-shaped pulses for all $g$ parameters, see Appendix B for details.. We label the latter unitary as $(\Pi/2)_{01}^{(L)}$. We use similar notation for qubit $R$. The vacuum state of the photonic lattice is |$\ket{\varnothing}_{\gamma}$. The topological transfer of a photonic excitation from the leftmost cavity to the rightmost through the edge states is $T_{LR}$.
@@
|$\ket{\Phi}_{LR}=\frac{1}{\sqrt{2}}\left[\ket{10}+e^{i\theta}\ket{01}\right],$ (6)
@@
$e^{i\theta}=[(-1)^{\ell/2+1}i]^{n_{d}}.$ (7)
See Appendix C for a detailed derivation of this relative phase [^2] This formula is related to those given in [19] for the SSH chain, where $\ell$ was defined differently and $n_{d}=n_{w}+1$., and Appendix D for the intermediate steps in this and all protocols.
3.2 Performance against disorder: concurrence
While the calculations reported so far demonstrate that this system generates entanglement in the ideal case, we need to explore the impact of disorder if we are to consider the system a scalable quantum platform. First of all, the final state will no longer be Eq. (6). Moreover, the final state has a nonzero probability amplitude of having a leftover photon in the photonic lattice, leading to decoherence in the state of the qubits. In this case, entanglement certification is not as simple as inspecting the state in Eq. (6), so instead we turn to concurrence, an entanglement measure [58], to perform this task. We remind the reader that, if a two-qubit state has high concurrence, any latter state with lower concurrence may be prepared from the former with only local operations and classical communication (LOCC); in particular, without using the photonic lattice. In this sense, Bell states have maximal concurrence. We now turn our attention to the disorder model used, and then show that the protocol can produce high-concurrence states despite disorder.